OpenStudy (anonymous):

find the limit limit as x approaches infinity of (7/n) + 1

OpenStudy (anonymous):

n approaches to infinity, right?

OpenStudy (matheducatormcg):

what happens as you divide 7 by a bigger, bigger, bigger number and then add 1? Try dividing 7 by 5000 and then add 1. What value are you getting close to?

OpenStudy (anonymous):

yes sorry n

OpenStudy (anonymous):

infinity?

OpenStudy (anonymous):

7/infinity =0 (as the larger the denominator, the smaller the value of the fraction is) so limit as n approaches infinity of (7/n) + 1 =1

OpenStudy (matheducatormcg):

\[\frac{7}{5000}+1=1.0014\] and well, 5000 is much smaller than infinity, so as you grow the denominator, the amount of zeros will grow in the decimal taking you closer to 1.

OpenStudy (anonymous):

so 1?

OpenStudy (matheducatormcg):

yes

OpenStudy (anonymous):

THANKS!

OpenStudy (anonymous):

what about (7+n)/n

OpenStudy (anonymous):

Its the same..

OpenStudy (anonymous):

it is?

OpenStudy (anonymous):

(7+n)/n = 7/n +n/n = 7/n +1 this is the same as the previous question

OpenStudy (anonymous):

sorry i meant (7+n)/n^2

OpenStudy (anonymous):

answer is 0

OpenStudy (anonymous):

and (7+n^2)/n

OpenStudy (anonymous):

the answer is not defined...

OpenStudy (anonymous):

the choices are : a)0 b)1 c)postive infinty

OpenStudy (matheducatormcg):

positive infinity

OpenStudy (matheducatormcg):

\[\lim_{n \rightarrow \infty} \frac{7+n^2}{n}=\lim_{n \rightarrow \infty} \frac{7}{n}+\frac{n^2}{n}\]

OpenStudy (matheducatormcg):

the 7/n goes to 0 and the n^2/n goes to infinity. so answer is infinity

OpenStudy (anonymous):

MUCH THANKS!!!!!